A) fond b) Find In FOX) = 20 T(OSX 530) sufficient Stat, MLE, MULLE, and mom of o 95% CI for o
Answer #1

Let X1, X2,..., Xn be random sample from linform (0,20). lle pdf of x as F(X)= t io <x<20 The likelihood function of Me... InMVUE: f(x) = n(x) i osx.120 n 2014 : 022526 - (200 200 E[] = I x. f(x) die m 20 o men20 ntl nt nt an” (2001) Enti 120 W V HenMLE: The likelihood function for o, Vegna ; 0<x5.5%,<20 The likelihood fention will bee maximized Tf. 200 > Xos, or 07 Hence,Les Y= x - xn=0> y=0 xn=203 y=1 an dX(W) = 20 Hence. The pelf of yo fY) = f(x) dxb f(y) =ny i osy<1Lex c7o such that, 1-9. Sfmody > 1-8. = 1.ch fence, PC c< Y<1) = 1-4 3 P(xin L < 1) = 1-4> PCB < < <a y) = 1-4 Hence, 1001-43% confidence Inlaval y & Xe Given, X= 0.05 then 95% confidence rulesual foro, g Z. (0.05)